OpenStudy (anonymous):

Solve the equation x + 2 = 22 for x. Answer 20 48 10 40

OpenStudy (anonymous):

20

OpenStudy (saifoo.khan):

subtract 2 from both sides.

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

its 1/2 x + 2 = 22 for x.

OpenStudy (anonymous):

I messed it up sorry I didnt put the right equation in

OpenStudy (anonymous):

then, 1/2x=20 and that is x=40

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

no problem

OpenStudy (saifoo.khan):

\[\frac 12 x + 2 = 22\]Subtract 2 from both sides, \[\frac12x = 20\]Multiply 2 from both sides, \[x=20 \times 2\]\[x=40\]

OpenStudy (anonymous):

great answer